John is 4 yrs older than Mary and Ainy is 2 yrs younger than Mary. If the sum of their ages is 41,how old will John be in 8 years' time?
Answer by Alan3354(69443)

John is 4 yrs older than Mary and Ainy is 2 yrs younger than Mary. If the sum of their ages is 41,how old will John be in 8 years' time?
============
J = M + 4
A = M - 2
J + M + A = 41
------------
Sub for J and A
M+4 + M + M-2 = 41
3M + 2 = 41
3M = 39
M = 13 (Mary's age now)
etc